OpenStudy (anonymous):

The sum of two numbers 51. The smaller number is 9 less than the larger number. What are the numbers?

OpenStudy (anonymous):

x-y=9 and x+y=51 => x=y+9 and y+9+y=51 => 2y=42 =>y=21

OpenStudy (anonymous):

and x=30

OpenStudy (anonymous):

lets say the smaller number is x then larger number should be x+9 sum up these: 2x+9=51 when we solve for x, x=21